# Mixed Hodge structures for vanishing cycles and orbifold cohomology

### v1

@@ 1 Introduction
Sabbah’s mixed Hodge structures have some interesting applications, which justify the interest shown in it: see for instance [8], [12], [13], [14], [24] for the relations with Frobenius manifolds and mirror symmetry. By [28, Theorem 4.3], it is also isomorphic to the limit mixed Hodge structure constructed by Steenbrink and Zucker in [39] on $H^{\bullet}(U,f^{-1}(t),\mathbb{Q})$ when $t\rightarrow\infty$.
@@ 2.1 The graded Milnor ring
Let $f(x)=\sum_{m}a_{m}x^{m}$, where $m=(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$ and $x^{m}=x_{1}^{m_{1}}\ldots x_{n}^{m_{n}}$, be a Laurent polynomial defined on $(\mathbb{C}^{*})^{n}$. Let $\operatorname{supp}(f)=\{m,a_{m}\neq 0\}$. The Newton polytope $P$ of $f$ is the convex hull of $\operatorname{supp}(f)$ in $\mathbb{R}^{n}$. We will say that $f$ is convenient if $P$ contains the origin as an interior point and that it is nondegenerate if, for each face $\Delta$ of $P$,
$x_{1}\frac{\partial f_{\Delta}}{\partial x_{1}},\ldots,x_{n}\frac{\partial f_{\Delta}}{\partial x_{n}}$
do not vanish simultaneously on $(\mathbb{C}^{*})^{n}$, where $f_{\Delta}(x):=\sum_{m\in\Delta}a_{m}x^{m}$.
@@ 2.3 Sabbah’s mixed Hodge structure
In this section, we recall Sabbah’s setting [28], [29]. Let $f$ be a convenient and nondegenerate Laurent polynomial on $(\mathbb{C}^{*})^{n}$, where $n\geq 2$. We will denote by $G$ the localized Laplace transform of its Gauss-Manin system and by $G_{0}$ its Brieskorn module:
@@
the latter being a finite dimensional vector space of dimension $\mu$, the global Milnor number of $f$. Actually, $G_{0}$ is a lattice in $G$, i.e $G_{0}$ is free over $\mathbb{C}[\theta]$ and $\mathbb{C}[\theta,\theta^{-1}]\otimes_{\mathbb{C}[\theta]}G_{0}=G$: this is shown in [13, Section 4], using the main properties of convenient and nondegenerate functions established in [22]. More generally, this is a characteristic property of tame functions, for which Sabbah’s mixed Hodge structures can be constructed [29, Section 13].
Let $\tau:=\theta^{-1}$. Then $G$ is a free $\mathbb{C}[\tau,\tau^{-1}]$-module, equipped with a derivation $\partial_{\tau}$ (induced from the formula $\partial_{\tau}[\omega]=[-f\omega]$ if $\omega\in\Omega^{n}(U)$ where the brackets denote the class in $G$). Let $V_{\bullet}G$ be the Kashiwara-Malgrange filtration of $G$ along $\tau=0$ as in [29, Section 1]. This is an increasing filtration, indexed by $\mathbb{Q}$, which satisfies the following properties:
@@
Sabbah’s mixed Hodge structures. By [29, Theorem 13.1], the triple $\operatorname{MHS}_{f}:=(H,F^{\bullet},W_{\bullet})$ is a mixed Hodge structure.
@@
where $[N]$ is induced by $N]:\operatorname{gr}_{F}^{i}H\rightarrow\operatorname{gr}_{F}^{i-1}H$: in particular, $N$ and $[N]$ have the same Jordan normal form (see [34, Theorem 7.1]).
@@ 2.4 Isomorphisms
Proof.  Note first that the assumptions of Proposition 2.2 imply that $f$ is convenient and nondegenerate. It remains to identify the graded ring appearing in [2, Theorem 1.1] with $A^{*}$. Recall that if $\partial:\mathcal{C}\rightarrow\mathcal{A}$ is a strict morphism of filtered module (i.e $\partial(\mathcal{C})\cap\mathcal{A}_{q}=\partial(\mathcal{C}_{q})$) we have $\operatorname{gr}(\mathcal{A}/\partial\mathcal{C})\cong\operatorname{gr}(\mathcal{A})/\operatorname{gr}(\partial)\operatorname{gr}(\mathcal{C})$, see for instance [22, Lemme 4.4]. We get the expected identification by applying this observation to the morphism
@@
$\partial(g_{1},\ldots,g_{n})=g_{1}x_{1}\frac{\partial f}{\partial x_{1}}+\ldots+g_{n}x_{n}\frac{\partial f}{\partial x_{n}},$
which is strict for the following choice of the filtration in $\mathcal{A}^{n}$: the Newton weight of $(0,\ldots,g_{i},\ldots,0)$ in is equal to $\nu(g_{i})-1$, see [22, Theorem 4.1].   $\Box$
@@
Proof.  We follow [34, Theorem 7.1]. For $\beta\in]0,1[$, we have the commutative diagram
@@ 3.2 Conjugations
Let us note that conjugation (25) appears implicitly in [42, Proposition 4.7], but it does not match with the conjugations appearing in Sabbah’s mixed Hodge structures, for which $2i\pi L_{\omega}(=2i\pi N)$ is real.
@@ 3.4 Hodge numbers
By Proposition 3.5, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ and $\operatorname{MHS}_{\operatorname{orb}}$ are the same. Therefore, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ satisfy the equalities of Proposition 3.12 and the conclusion of its corollaries. A formula for the Hodge numbers of the mixed Hodge structure $\operatorname{MHS}_{f}$ is also given in [42, Corollary 4.10] in terms of the $h$-vector of the Newton polytope $P$ of $f$.
@@ 4 Interlude: the Hodge-Tate condition
Proof.  The equivalence between the first and the third point can be found in [7, Proposition 2.1.5], and the equivalence between the first and the last point is precisely [30, Corollary 2.3]. The remaining assertion is straightforward.   $\Box$
@@ 5.1.1 Orbifold Poincaré-Saito duality
II denoting the involution $I:X_{\ell}\rightarrow X_{\ell^{-1}}$ alluded to in Section 2.2. We refer to [3] or [23, III 3.3] for the definition of the orbifold integral. We will call $g$ the Poincaré-Saito duality.
@@ 5.1.2 A polarization
by [23, Proposition IV.3.13] (this is the volume of the polar polytope $P^{*}$ of $P$; recall that $\mu=q_{0}+q_{1}+\ldots+q_{n}$). Since the twisted sectors $\mathcal{X}_{\ell}$ are also weighted projective spaces, we have analogous formulas for $\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}N^{n_{\ell}}$: for $\ell\in F$ we have
@@ Appendix A Appendix (polarized mixed-Hodge structures)
Let $N:V\rightarrow V$ be a linear map on a finite dimensional vector space over a field of characteristic 0, which satisfies $N^{r+1}=0$. Then (see [35, Lemma 6.4]),
@@ B.1 Foreword: about the duality of $\mathcal{D}$-modules
Given a $\mathbb{C}[\tau,\tau^{-1}]$-module $G$, we denote by $\overline{G}$ the $\mathbb{C}$-vector space $G$ equipped with the new module structure $p(\tau).g:=p(-\tau)g$. We will denote by $\overline{g}$ the elements of $\overline{G}$. If moreover $G$ is equipped with a connection $\partial_{\tau}$ then so is $G$ and we have $\partial_{\tau}\overline{g}:=\overline{-\partial_{\tau}g}$. When $G$ is the Fourier-transform of the Gauss-Manin system of $f$ as in Section 2.3, duality of $\mathcal{D}$-modules gives (see [29, Section 3] for details and also [32, Section 2.7] for the local model) the existence of nondegenerate $\mathbb{C}[\tau,\tau^{-1}]$-bilinear pairings
@@ B.2 Application to weighted projective spaces
where $q^{q}=q_{0}^{q_{0}}\ldots q_{n}^{q_{n}}$ and $c_{k,\ell}$ is a positive rational number expressed with the help of the weights $q_{0},\ldots,q_{n}$, see for instance [23, Lemme VI.1.18]. Moreover, $S(\omega_{k},\overline{\omega_{\ell}})\in\mathbb{C}\tau^{-n}$ and we will write

### v2

@@ 1 Introduction
Sabbah’s mixed Hodge structures have some interesting applications, which justify the interest shown in it: see for instance [8], [12], [13], [14], [25] for the relations with Frobenius manifolds and mirror symmetry. By [29, Theorem 4.3], it is also isomorphic to the limit mixed Hodge structure constructed by Steenbrink and Zucker in [40] on $H^{\bullet}(U,f^{-1}(t),\mathbb{Q})$ when $t\rightarrow\infty$ (we set here $U:=(\mathbb{C}^{*})^{n}$).
@@ 2.1 The graded Milnor ring
Let $f(u)=\sum_{m}a_{m}u^{m}$, where $m=(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$ and $u^{m}=u_{1}^{m_{1}}\ldots u_{n}^{m_{n}}$, be a Laurent polynomial defined on $(\mathbb{C}^{*})^{n}$. Let $\operatorname{supp}(f)=\{m,a_{m}\neq 0\}$. The Newton polytope $P$ of $f$ is the convex hull of $\operatorname{supp}(f)$ in $\mathbb{R}^{n}$. We will say that $f$ is convenient if $P$ contains the origin as an interior point and that it is nondegenerate if, for each face $\Delta$ of $P$,
$u_{1}\frac{\partial f_{\Delta}}{\partial u_{1}},\ldots,u_{n}\frac{\partial f_{\Delta}}{\partial u_{n}}$
do not vanish simultaneously on $(\mathbb{C}^{*})^{n}$, where $f_{\Delta}(u):=\sum_{m\in\Delta}a_{m}u^{m}$.
@@ 2.3 Sabbah’s mixed Hodge structure
In this section, we recall Sabbah’s setting [29], [30]. Let $f$ be a convenient and nondegenerate Laurent polynomial on $U:=(\mathbb{C}^{*})^{n}$, where $n\geq 2$. We will denote by $G$ the localized Laplace transform of its Gauss-Manin system and by $G_{0}$ its Brieskorn module:
@@
the latter being a finite dimensional vector space of dimension $\mu$, the global Milnor number of $f$. Actually, $G_{0}$ is a lattice in $G$, i.e $G_{0}$ is free over $\mathbb{C}[\theta]$ and $\mathbb{C}[\theta,\theta^{-1}]\otimes_{\mathbb{C}[\theta]}G_{0}=G$: this is shown in [13, Section 4], using the main properties of convenient and nondegenerate functions established in [23]. More generally, this is a characteristic property of tame functions, for which Sabbah’s mixed Hodge structures can be constructed [30, Section 13].
Let $\tau:=\theta^{-1}$. Then $G$ is a free $\mathbb{C}[\tau,\tau^{-1}]$-module, equipped with a derivation $\partial_{\tau}$ (induced from the formula $\partial_{\tau}[\omega]=[-f\omega]$ if $\omega\in\Omega^{n}(U)$ where the brackets denote the class in $G$). Let $V_{\bullet}G$ be the Kashiwara-Malgrange filtration of $G$ along $\tau=0$ as in [30, Section 1]. This is an increasing filtration, indexed by $\mathbb{Q}$, which satisfies the following properties:
@@
Sabbah’s mixed Hodge structures. By [30, Theorem 13.1], the triple $\operatorname{MHS}_{f}:=(H,F^{\bullet},W_{\bullet})$ is a mixed Hodge structure.
@@
where $[N]$ is induced by $N]:\operatorname{gr}_{F}^{i}H\rightarrow\operatorname{gr}_{F}^{i-1}H$: in particular, $N$ and $[N]$ have the same Jordan normal form (see [35, Theorem 7.1]).
@@ 2.4 Isomorphisms
Proof.  Note first that the assumptions of Proposition 2.2 imply that $f$ is convenient and nondegenerate. It remains to identify the graded ring appearing in [2, Theorem 1.1] with $A^{*}$. Recall that if $\partial:\mathcal{C}\rightarrow\mathcal{A}$ is a strict morphism of filtered module (i.e $\partial(\mathcal{C})\cap\mathcal{A}_{q}=\partial(\mathcal{C}_{q})$) we have $\operatorname{gr}(\mathcal{A}/\partial\mathcal{C})\cong\operatorname{gr}(\mathcal{A})/\operatorname{gr}(\partial)\operatorname{gr}(\mathcal{C})$, see for instance [23, Lemme 4.4]. We get the expected identification by applying this observation to the morphism
@@
$\partial(g_{1},\ldots,g_{n})=g_{1}u_{1}\frac{\partial f}{\partial u_{1}}+\ldots+g_{n}u_{n}\frac{\partial f}{\partial u_{n}},$
which is strict for the following choice of the filtration in $\mathcal{A}^{n}$: the Newton weight of $(0,\ldots,g_{i},\ldots,0)$ in is equal to $\nu(g_{i})-1$, see [23, Theorem 4.1].   $\Box$
@@
Proof.  We follow [35, Theorem 7.1]. For $\beta\in]0,1[$, we have the commutative diagram
@@ 3.2 Conjugations
Let us note that conjugation (25) appears implicitly in [43, Proposition 4.7], but it does not match with the conjugations appearing in Sabbah’s mixed Hodge structures, for which $2i\pi L_{\omega}(=2i\pi N)$ is real.
@@ 3.4 Hodge numbers
By Proposition 3.5, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ and $\operatorname{MHS}_{\operatorname{orb}}$ are the same. Therefore, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ satisfy the equalities of Proposition 3.12 and the conclusion of its corollaries.
Remark 3.16
A formula for the Hodge numbers of the mixed Hodge structure $\operatorname{MHS}_{f}$ is also given in [43, Corollary 4.10] in terms of the $h$-vector of the Newton polytope $P$ of $f$. As already emphasized in the introduction, such Hodge numbers are also computed using different techniques in other papers (see for instance [21], [26], [41]). The computation (via mirror symmetry) of Hodge numbers of Calabi-Yau varieties produced by crepant resolutions of toric varieties coming from reflexive polytopes is discussed in [18].
@@ 4 Interlude: the Hodge-Tate condition
Proof.  The equivalence between the first and the third point can be found in [7, Proposition 2.1.5], and the equivalence between the first and the last point is precisely [31, Corollary 2.3]. The remaining assertion is straightforward.   $\Box$
@@ 5.1.1 Orbifold Poincaré-Saito duality
II denoting the involution $I:X_{\ell}\rightarrow X_{\ell^{-1}}$ alluded to in Section 2.2. We refer to [3] or [24, III 3.3] for the definition of the orbifold integral. We will call $g$ the Poincaré-Saito duality.
@@ 5.1.2 A polarization
by [24, Proposition IV.3.13] (this is the volume of the polar polytope $P^{*}$ of $P$; recall that $\mu=q_{0}+q_{1}+\ldots+q_{n}$). Since the twisted sectors $\mathcal{X}_{\ell}$ are also weighted projective spaces, we have analogous formulas for $\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}N^{n_{\ell}}$: for $\ell\in F$ we have
@@ Appendix A Appendix (polarized mixed-Hodge structures)
Let $N:V\rightarrow V$ be a linear map on a finite dimensional vector space over a field of characteristic 0, which satisfies $N^{r+1}=0$. Then (see [36, Lemma 6.4]),
@@ B.1 Foreword: about the duality of $\mathcal{D}$-modules
Given a $\mathbb{C}[\tau,\tau^{-1}]$-module $G$, we denote by $\overline{G}$ the $\mathbb{C}$-vector space $G$ equipped with the new module structure $p(\tau).g:=p(-\tau)g$. We will denote by $\overline{g}$ the elements of $\overline{G}$. If moreover $G$ is equipped with a connection $\partial_{\tau}$ then so is $G$ and we have $\partial_{\tau}\overline{g}:=\overline{-\partial_{\tau}g}$. When $G$ is the Fourier-transform of the Gauss-Manin system of $f$ as in Section 2.3, duality of $\mathcal{D}$-modules gives (see [30, Section 3] for details and also [33, Section 2.7] for the local model) the existence of nondegenerate $\mathbb{C}[\tau,\tau^{-1}]$-bilinear pairings
@@ B.2 Application to weighted projective spaces
where $q^{q}=q_{0}^{q_{0}}\ldots q_{n}^{q_{n}}$ and $c_{k,\ell}$ is a positive rational number expressed with the help of the weights $q_{0},\ldots,q_{n}$, see for instance [24, Lemme VI.1.18]. Moreover, $S(\omega_{k},\overline{\omega_{\ell}})\in\mathbb{C}\tau^{-n}$ and we will write